OpenStudy (lgbasallote):

The position pf a particle moving along the x-axis is given by x = (21 + 22t - 6t^2)m, where t is in seconds. What is the average velocity during the time interval t = 1 s to t = 3 s

OpenStudy (lgbasallote):

i know \[f' (t) = 22 - 12t\]

OpenStudy (lgbasallote):

so do i do \[22 - 12(3) - [22 - 12(1)]\]

OpenStudy (lgbasallote):

oh wait...i know what i did wrong...hold on

OpenStudy (lgbasallote):

i do \[22 - 12(3-1)\] right?

OpenStudy (lgbasallote):

-2?

OpenStudy (lgbasallote):

yup. it's -2.

OpenStudy (lgbasallote):

thanks for watching @curiousshubham

OpenStudy (lgbasallote):

but seriously...is my solution right?

OpenStudy (anonymous):

hey sorry i was outside........ let me check...

OpenStudy (lgbasallote):

my solution is \[22 - 12(3-1)\] \[\implies - 2\]

OpenStudy (lgbasallote):

\[f(t) = 21 + 22t - 6t^2\] \[\implies f'(t) = 22 - 12t\] \[\implies 22 - 12(3-1)\] \[\implies - 2\]

OpenStudy (anonymous):

hey its right.. but wait em still confused but don't know why...

OpenStudy (anonymous):

yeah right....

OpenStudy (lgbasallote):

yeah right it's right?

OpenStudy (lgbasallote):

i'll take that as a yes

OpenStudy (anonymous):